OpenStudy (anonymous):

What does it mean to check if two vectors, say Cos[Pix] and Sin[Pix], are linear independent in C[0,1]?

OpenStudy (anonymous):

Deos that mean when x=0 and 1?

OpenStudy (amistre64):

i believe C is the complex plane

OpenStudy (anonymous):

So C[0,1] is simple saying no imaginary numbers?

OpenStudy (amistre64):

if vectors are independant, then they are not scalar multiples of each other, nor can they be created from a combination of other vectors in the set

OpenStudy (amistre64):

the C[0,1] is confusing without some context. What does this notation represent in your tesxt?

OpenStudy (anonymous):

I know they are linear independent because the wronskain determinant shows them to be Pi but say you have a Wronskian determinant for two vectors with a non-zero solution and an X.

OpenStudy (anonymous):

This is exactly written from text: For each of the following. show that the given vecotrs are linearly independent in C[0,1]: it then lists examples a, b, c, and d.

OpenStudy (anonymous):

W[x^3/2, x^5/2]=5/2x^2-3/2x^3

OpenStudy (anonymous):

That above answer shows it to be linearly independent but only for certain x values. I figuredd C[0,1] had something to do with checking certain x values

OpenStudy (anonymous):

whatcha think? I think it has something to do with the vector space or the constants c1, c2

OpenStudy (amistre64):

still trying to determine what "in C[0,1]" is refering to. Any ideas phi?

OpenStudy (anonymous):

must have to do with x [0,1]

OpenStudy (amistre64):

\[W\begin{vmatrix} cos(\pi~x)&sin(\pi~x)\\ -\pi~sin(\pi~x)&\pi~cos(\pi~x) \end{vmatrix}=\pi~(cos^2(\pi~x)+sin^2(\pi~x)=\pi\]

OpenStudy (anonymous):

Okay I found an example showing C[-infinity,infinity] for W[1,x,x^2,x^3]=12 making it linearly independent for all x. So example above is linearly independent for C[0,1] and all x.

OpenStudy (anonymous):

it's strange

OpenStudy (amistre64):

i take it this is not a first year linear algebra question :) I keep reading about the wronskian being applied of an interval [a,b] and it relating to a differential equation setup

OpenStudy (amistre64):

Let \(f_1\) and \(f_2\) be solutions to the diffyQ: \(y''+p(t)y'+q(t)=0\). Then W(\(f_1,f_2\)) is either zero for all t, or never zero

OpenStudy (anonymous):

No, it is actually. I did a google search before coming here and only found diff. eq stuff but we haven't gone into any of that.

OpenStudy (anonymous):

Yeah i was on that page earlier

OpenStudy (amistre64):

well, if its just first year stuff, then Two vectors are linearly independent if \(c_1x+c_2y=0\) only if c1=c2=0

OpenStudy (amistre64):

and i believe the C[0,1] is stating, within the interval from 0 to 1

OpenStudy (amistre64):

\[c_1~cos(x~\pi)+c_2~sin(x~\pi)=0\] http://www.wolframalpha.com/input/?i=cos%28pi+x%29%2Bsin%28pi+x%29%3D0 it appears that if we let c1=c2=1, that we get cos(pi x) + sin(pi x) = 0 between the interval 0 to 1

OpenStudy (anonymous):

I see but if we are assuming C[0,1] has to do with c1, and c2 then the interval [0,1] for x no longer is what they are asking for?...correct my if im wrong.

OpenStudy (amistre64):

C[0,1] im reading as an interval from x=0 to x=1 or simply the values of cos(t), and sin(t) from 0 to pi along that interval we see that sin(t) = -cos(t) , when t = 3pi/4 . Since along the interval, we encounter an instance such that the one vector is a scalar multiple of the other ... id say that they cannot be classified as independant. But thats real fuzzy since i took this stuff awhile ago

OpenStudy (amistre64):

i dont have my textbook handy to reference tho

OpenStudy (anonymous):

i see, for now I will stick with the wronskian determinant that says they are independent but I like the reasoning. Thanks for your time! If I get any answers I will post them soon. Have a good day! (:

OpenStudy (amistre64):

good luck ;) and ill most likely have time to review this more over the weekend to see if my memory has failed me lol

OpenStudy (anonymous):

thanks amistre64